Use the 5-step problem-solving process. The sum of 3 consecutive integers is 423. Find the integers.
Answer by nyc_function(2741)

Let x = first integer
Let x + 1 = second integer
Let x + 2 = third integer
x + x + 1 + x + 2 = 423
3x + 3 = 423
3x = 423 - 3

3x = 420

x = 420/3

x = 140

Now that we know x = 140, we can find all three integers.

First integer = x or 140

Second integer = x + 1 or 141

Third integer = x + 2 or 142

Understand?